(D2+DD')Z=sin(x+y)



1
Expert's answer
2022-02-24T11:51:11-0500

The auxiliary equation is "m^2+m=0."

The roots of the equation are "m(m+1)=0\implies m =0,-1".

The complementary function is "C.F = f_1(y)+f_2(y-x)."

The particular integral is

"\begin{aligned}
P.I &= \dfrac{1}{D^2+DD'}\sin(x+y)\\
&=\dfrac{1}{-1-1}\sin(x+y)~~(\text{Replacing $D^2=-1^2, DD'=-1$})\\
&=-\dfrac{1}{2}\sin(x+y)
\end{aligned}"


The solution is "z = f_1(y)+f_2(y-x)-\dfrac{\sin(x+y)}{2}"